# Coupler Position Optimization and Channel Estimation for Flexible Coupler Aided Multiuser Communication

### v1

@@ Abstract
In this paper, we propose a distributed flexible coupler (FC) array to enhance communication performance with low hardware cost. At each FC antenna, there is one fixed-position active antenna and multiple passive couplers that can move within a designated region around the active antenna. Moreover, each FC antenna is equipped with a local processing unit (LPU). All LPUs exchange signals with a central processing unit (CPU) for joint signal processing. We study an FC-aided multiuser multiple-input multiple-output (MIMO) system, where an FC array base station (BS) is deployed to enhance the downlink communication between the BS and multiple single-antenna users. We formulate optimization problems to maximize the achievable sum rate of users by jointly optimizing the coupler positions and digital beamforming, subject to movement constraints on the coupler positions and the transmit power constraint. To address the resulting nonconvex optimization problem, the digital beamforming is expressed as a function of the FC position vectors, which are then optimized using the proposed distributed coupler position optimization algorithm. Considering a structured time domain pattern of pilots and coupler positions, pilot-assisted centralized and distributed channel estimation algorithms are designed under the FC array architecture. Simulation results demonstrate that the distributed FC array achieves substantial rate gains over conventional benchmarks in multiuser systems without moving active antennas, and approaches the performance of fully active arrays while significantly reducing hardware cost and power consumption. Moreover, the proposed centralized channel estimation algorithms outperform the benchmark schemes in terms of both pilot overhead and channel reconstruction accuracy. It is verified that reconfiguring passive coupler positions offers a cost-effective way to enhance wireless communication performance without moving active antennas.
@@ I Introduction
To overcome the above limitations, we propose in this paper a new distributed FC array, as illustrated in Fig. 1, where each FC antenna is equipped with a local processing unit (LPU) to enable decentralized and parallel processing, thereby alleviating the computational burden at the CPU while improving communication performance. Moreover, by dispensing with the requirement that each antenna be connected to a dedicated RF chain, the FC array consists of multiple FC antennas, each comprising a single active antenna connected to one RF chain and multiple passive couplers. The passive couplers can be independently adjusted in position around their associated active antenna, while the active antenna remains fixed at the transceiver. By exploiting coupler movement and the mutual coupling between neighboring antenna and coupler elements, the passive couplers radiate through excitation induced by their associated active antenna. In the proposed distributed FC array, all LPUs operate in a decentralized and parallel manner and exchange signals with the CPU for joint signal processing, such as channel estimation, precoding, and combining. The proposed architecture leverages both the CPU and the LPUs to not only offload computational tasks from the CPU but also reduce the frequency of data exchange between the distributed FC antennas and the CPU by exploiting local computation and baseband signal processing at the LPUs. As a result, the overall processing cost and latency are effectively reduced.
Different from six-dimensional movable antennas (6DMA) [21, 26, 25, 22], which enhance communication performance by translating and/or rotating active antennas together with their RF chains over relatively large regions[20, 17, 40, 44, 16, 4, 30, 23, 8, 37, 33, 39, 31, 28, 27], in an FC system, only the passive couplers are repositioned within a small local region where electromagnetic coupling is strong, while the active antennas and their RF chains remain fixed. This design substantially reduces transmit power consumption and mechanical control complexity, and enables faster and more practical position reconfiguration. Thus, the distributed FC array features advantages such as a compact structure, low profile, and lightweight, which make it suitable for small-sized terminals and practical deployment.
@@
  • We propose a distributed FC array to enhance communication performance and reduce centralized processing complexity by equipping each FC antenna with a local LPU. The FC array consists of multiple FC antennas, each containing one fixed-position active antenna and multiple passive couplers that can move within a designated region around each active antenna to enhance the multiuser communication. The distributed FC array features a compact structure without moving the active antennas and their RF chains, making it particularly attractive for large-scale antenna deployments in small-sized devices.
@@
  • Based on the distributed FC array, we formulate an optimization problem to maximize the achievable sum rate of users by determining the positions of the passive couplers and the transmit beamforming, subject to their practical movement and transmit power constraints. To solve this nonconvex problem, we develop a successive convex approximation (SCA)-based distributed coupler-position optimization algorithm, where the LPUs update the coupler positions in parallel via local projections and the CPU broadcasts the required gradient information and updates the minimum mean square error (MMSE) precoder.
@@
  • By exploiting a structured training protocol with multiple pilot blocks and blockwise fixed coupler positions, we develop two pilot-assisted channel estimation algorithms under the distributed FC architecture, namely, centralized scheme and distributed scheme. In the centralized scheme, the CPU stacks pilot-correlated observations across all FC antennas and performs sparse recovery to estimate the angular support and path gains, enabling channel reconstruction for arbitrary coupler positions. In the distributed scheme, each LPU conveys low-overhead support information to the CPU, which carries out lightweight fusion and path gain estimation using low-dimensional statistics.
@@
  • Extensive simulations validate the effectiveness of the proposed distributed FC array and the associated optimization and channel estimation designs. The results show that the proposed FC array can closely approach the achievable rate of a fully active array while using significantly fewer active antennas, due to the interference mitigation, mechanical beamforming gain, and fading mitigation enabled by coupler repositioning. Moreover, simulations demonstrate the efficiency of the proposed channel estimation schemes.
Figure 1: Proposed distributed FC array for multiuser systems.
@@
The remainder of this paper is organized as follows. Section II describes the system model, including the distributed FC array structure, the channel model, mechanical beamforming, and the signal model of the distributed FC array. Section III formulates the optimization problem for maximizing the achievable sum rate through coupler position design and then presents the proposed distributed coupler position optimization algorithm. Section IV presents the proposed centralized and distributed channel estimation algorithms for the designed FC array. Section V provides numerical results and corresponding discussions. Finally, Section VI concludes this work.
@@ II-A Distributed Flexible Coupler Array
As shown in Fig. 1, the proposed FC array consists of $M$ FC antennas. Each FC antenna comprises one fixed-position active antenna and $N$ passive couplers, which can move within a designated region, i.e., $\mathcal{C}_{m}$ surrounding the $m$-th active antenna. Each active antenna is connected to a single RF chain. The positions of passive couplers are mechanically adjustable with the aid of actuation components such as micro-electro-mechanical systems (MEMS), which offer low power consumption and fast response. Moreover, each FC antenna is equipped with local computing hardware, i.e., an LPU, which carries out the necessary baseband processing tasks, such as channel estimation and signal precoding and combining, in a decentralized and parallel fashion, as shown in Fig. 1. Each LPU is connected to the common CPU at the base station (BS) via separate flexible cables, such as coaxial cables. These cables supply power to the FC antenna and also facilitate control and signal exchange between the LPU and the CPU. The users are each assumed to be equipped with a single fixed-position antenna. In this architecture, the signals radiated by the passive couplers are generated through induced currents excited by their associated active antenna via near-field electromagnetic coupling, without requiring additional RF chains or feeding networks. By reconfiguring only the passive coupler positions within a small local region, the proposed FC array enhances wireless communication performance while avoiding the movement of active antennas or RF chains, thereby significantly reducing mechanical control complexity and energy consumption.
@@ II-B Channel Model
As illustrated in Fig. 1, we consider a downlink multiuser wireless communication system where a FC array is deployed at the BS. The BS serves $K$ single-antenna users. The FC array can be reconfigured by mechanically adjusting the positions of its passive couplers.
@@
where $\mathbf{p}_{m}=[\mathbf{p}_{1,m}^{\mathrm{T}},\mathbf{p}_{2,m}^{\mathrm{T}},\ldots,\mathbf{p}_{N,m}^{\mathrm{T}}]^{\mathrm{T}}\in\mathbb{R}^{2N}$ is the position vector of the couplers in the $m$-th FC antenna.
@@
where $\mathbf{a}_{\mathrm{C},k,\ell}(\mathbf{p}_{m})$ denotes the coupler steering vector associated with user $k$, path $\ell$, and the $m$-th FC antenna, and is given by
@@
Based on (3) and (9), the overall multipath channel vector from the FC array to user $k$ is given by
@@
which dependence on coupler positions $\mathbf{p}$ while the active antenna positions are fixed. Collecting the channel vectors of all users yields the following multiuser MIMO channel matrix
@@ II-C Mechanical Beamforming
Let $\widetilde{\mathbf{Z}}\in\mathbb{C}^{M\times M}$ denote the mutual impedance among all active antennas and $\hat{\mathbf{Z}}(\mathbf{p})\in\mathbb{C}^{MN\times MN}$ denote the mutual impedance among all couplers. The overall mutual impedance matrix of the FC array is thus given by
@@
For practical implementation considerations, we assume that the active antennas are physically isolated by metallic plates. The center-to-center distance between any two active antennas is assumed to be significantly larger than the distance between each active antenna and its associated passive couplers, as well as the distances among the couplers themselves. Consequently, the mutual coupling between ports associated with different FC antennas becomes negligible and can be ignored. Under this assumption, the overall mutual impedance matrix exhibits a block-diagonal structure. Accordingly, we define the mutual impedance sub-matrix corresponding to the $m$-th FC antenna as
@@
Let $X_{n,m}$ denote the load impedance at the port of the $n$-th coupler associated with the $m$-th active antenna. The diagonal load-impedance matrix connected to the $N$ couplers of the $m$-th FC antenna is given by [9]
@@
Let $s\in\mathbb{C}$ denote the transmit information symbol with zero mean and unit variance. Antenna current vector $\mathbf{i}$ is adopted as the equivalent transmit precoding (beamforming) vector, such that the radiated signal is given by $\mathbf{i}s$. For the $m$-th FC antenna, define $\mathbf{v}_{\mathrm{C},m}\in\mathbb{C}^{N\times 1}$ as the port-voltage vector across the load impedances at the network side, and define the load-voltage drop along the direction of coupler current vector $\mathbf{i}_{\mathrm{C},m}\in\mathbb{C}^{N\times 1}$. By Kirchhoff’s voltage law [11], we have
@@
is the local precoding/beamforming vector of the $m$-th FC antenna. We refer to $\mathbf{w}_{m}(\mathbf{p}_{m})$ as the mechanical beamforming vector, since its beamforming weights can be reconfigured through adjusting the coupler positions, $\mathbf{p}_{m}$. Stacking the mechanical beamforming vectors of all FC antennas gives the following block-diagonal mechanical beamforming matrix,
@@ II-D Signal Model
For a given coupler-position vector $\mathbf{p}$, we use the active-antenna current vector as the equivalent digital beamforming vector across the $M$ FC antennas. Specifically, for user $k\in\mathcal{K}$, we define
@@ III-B MMSE-Based Digital Beamforming
For given coupler positions $\mathbf{p}$, the FC system reduces to a conventional downlink multiuser MIMO with precoder $\mathbf{F}$ under (48) and effective channel matrix
@@ III-C Coupler Position Optimization Algorithm
Problem (P2) is still nonconvex due to the complicated dependence of $R_{\rm MMSE}(\mathbf{p})$ on $\mathbf{p}$ and the nonconvex spacing constraints in (42d). However, the two constraints (42c) and (42d) are separable across different FC antennas, which enables a distributed optimization algorithm.
@@
denote the local feasible set of coupler positions at the $m$-th FC antenna. Due to the spacing constraints, $\mathcal{S}_{m}$ is in general nonconvex. To apply the SCA framework with convergence guarantees, we construct at each iteration a convex inner approximation of $\mathcal{S}_{m}$. Specifically, for each pair $(n,n^{\prime})$ with $n\neq n^{\prime}$, we define
@@
For problem (P2), the feasible set is separable across the $M$ FC antennas through $\{\mathcal{S}_{m}\}_{m=1}^{M}$, while the objective function $R_{\rm MMSE}(\mathbf{p})$ couples all antennas through the effective channels and the MMSE-based digital precoder. Let
@@
Algorithm 1 summarizes the distributed coupler position optimization procedure based on a distributed SCA scheme, where each FC antenna maximizes a strongly concave local surrogate of the global sum rate. Under standard smoothness assumptions on $J(\mathbf{p})$ and a diminishing step size sequence, the generated sequence $\{\mathbf{p}^{(t)}\}$ yields nondecreasing sum rate $J(\mathbf{p}^{(t)})$, and every limit point of $\{\mathbf{p}^{(t)}\}$ is a stationary solution of problem (P2). The per-iteration computational complexity is dominated by the computation of the MMSE-based precoder and the gradients, and is given by $\mathcal{O}(K^{2}M+K^{3}+KMN+MN^{2})$. Let $T_{\max}$ denote the number of iterations required by the stopping criterion in Algorithm 1. The overall complexity therefore scales as $\mathcal{O}\!\big(T_{\max}(K^{2}M+K^{3}+KMN+MN^{2})\big)$. At each iteration, the CPU broadcasts the local gradients to all LPUs, and each LPU uploads its updated coupler-position vector to the CPU. Hence, the communication overhead scales on the order of $MN$ scalar parameters.

### v2

# Coupler Position Optimization and Channel Estimation for Flexible Coupler Antenna Aided Multiuser Communication

@@ Abstract
In this paper, we propose a distributed flexible coupler antenna (FCA) array to enhance communication performance with low hardware cost. At each FCA, there is one fixed-position active antenna and multiple passive couplers that can move within a designated region around the active antenna. Moreover, each FCA is equipped with a local processing unit (LPU). All LPUs exchange signals with a central processing unit (CPU) for joint signal processing. We study an FCA-aided multiuser multiple-input multiple-output (MIMO) system, where an FCA array base station (BS) is deployed to enhance the downlink communication between the BS and multiple single-antenna users. We formulate optimization problems to maximize the achievable sum rate of users by jointly optimizing the coupler positions and digital beamforming, subject to movement constraints on the coupler positions and the transmit power constraint. To address the resulting nonconvex optimization problem, the digital beamforming is expressed as a function of the coupler position vectors, which are then optimized using the proposed distributed coupler position optimization algorithm. Considering a structured time domain pattern of pilots and coupler positions, pilot-assisted centralized and distributed channel estimation algorithms are designed under the FCA array architecture. Simulation results demonstrate that the distributed FCA array achieves substantial rate gains over conventional benchmarks in multiuser systems without moving active antennas, and approaches the performance of fully active arrays while significantly reducing hardware cost and power consumption. Moreover, the proposed centralized channel estimation algorithm outperforms the benchmark schemes in terms of both pilot overhead and channel reconstruction accuracy. It is verified that reconfiguring passive coupler positions offers a cost-effective way to enhance wireless communication performance without moving active antennas.
@@ I Introduction
To overcome the above limitations, we propose in this paper a new distributed flexible coupler antenna (FCA) array, as illustrated in Fig. 1, where each FCA is equipped with a local processing unit (LPU) to enable decentralized and parallel processing, thereby alleviating the computational burden at the CPU while improving communication performance. Moreover, by dispensing with the requirement that each antenna be connected to a dedicated RF chain, the FCA array consists of multiple FCAs, each comprising a single active antenna connected to one RF chain and multiple passive couplers. The passive couplers can be independently adjusted in position around their associated active antenna, while the active antenna remains fixed at the transceiver. By exploiting coupler movement and the mutual coupling between neighboring antenna and coupler elements, the passive couplers radiate through excitation induced by their associated active antenna. In the proposed distributed FCA array, all LPUs operate in a decentralized and parallel manner and exchange signals with the CPU for joint signal processing, such as channel estimation, precoding, and combining. The proposed architecture leverages both the CPU and the LPUs to not only offload computational tasks from the CPU but also reduce the frequency of data exchange between the distributed FCAs and the CPU by exploiting local computation and baseband signal processing at the LPUs. As a result, the overall processing cost and latency are effectively reduced.
Different from six-dimensional movable antennas (6DMA) [22, 27, 26, 23], which enhance communication performance by translating and/or rotating active antennas together with their RF chains over relatively large regions[21, 18, 41, 45, 17, 4, 31, 24, 8, 38, 34, 40, 32, 29, 28], in an FCA system, only the passive couplers are repositioned within a small local region where electromagnetic coupling is strong, while the active antennas and their RF chains remain fixed. This design substantially reduces transmit power consumption and mechanical control complexity, and enables faster and more practical position reconfiguration. Thus, the distributed FCA array features advantages such as a compact structure, low profile, and lightweight, which make it suitable for small-sized terminals and practical deployment.
@@
  • We propose a distributed FCA array to enhance communication performance and reduce centralized processing complexity by equipping each FCA with a local LPU. The FCA array consists of multiple FCAs, each containing one fixed-position active antenna and multiple passive couplers that can move within a designated region around each active antenna to enhance the multiuser communication. The distributed FCA array features a compact structure without moving the active antennas and their RF chains, making it particularly attractive for large-scale antenna deployments in small-sized devices.
@@
  • Based on the distributed FCA array, we formulate an optimization problem to maximize the achievable sum rate of users by determining the positions of the passive couplers and the transmit beamforming, subject to their practical movement and transmit power constraints. To solve this nonconvex problem, we develop a successive convex approximation (SCA)-based distributed coupler-position optimization algorithm, where the LPUs update the coupler positions in parallel via local projections and the CPU broadcasts the required gradient information and updates the minimum mean square error (MMSE) precoder.
@@
  • By exploiting a structured training protocol with multiple pilot blocks and blockwise fixed coupler positions, we develop two pilot-assisted channel estimation algorithms under the distributed FCA architecture, namely, centralized scheme and distributed scheme. In the centralized scheme, the CPU stacks pilot-correlated observations across all FCAs and performs sparse recovery to estimate the angular support and path gains, enabling channel reconstruction for arbitrary coupler positions. In the distributed scheme, each LPU conveys low-overhead support information to the CPU, which carries out lightweight fusion and path gain estimation using low-dimensional statistics.
@@
  • Extensive simulations validate the effectiveness of the proposed distributed FCA array and the associated optimization and channel estimation designs. The results show that the proposed FCA array can closely approach the achievable rate of a fully active array while using significantly fewer active antennas, due to the interference mitigation, mechanical beamforming gain, and fading mitigation enabled by coupler repositioning. Moreover, simulations demonstrate the efficiency of the proposed channel estimation schemes.
Figure 1: Proposed distributed FCA array for multiuser systems.
@@
The remainder of this paper is organized as follows. Section II describes the system model, including the distributed FCA array structure, the channel model, mechanical beamforming, and the signal model of the distributed FCA array. Section III formulates the optimization problem for maximizing the achievable sum rate through coupler position design and then presents the proposed distributed coupler position optimization algorithm. Section IV presents the proposed centralized and distributed channel estimation algorithms for the designed FCA array. Section V provides numerical results and corresponding discussions. Finally, Section VI concludes this work.
@@ II-A Distributed Flexible Coupler Array
As shown in Fig. 1, the proposed FCA array consists of $M$ FCAs. Each FCA comprises one fixed-position active antenna and $N$ passive couplers, which can move within a designated region, i.e., $\mathcal{C}_{m}$ surrounding the $m$-th active antenna. Each active antenna is connected to a single RF chain. The positions of passive couplers are mechanically adjustable with the aid of actuation components such as micro-electro-mechanical systems (MEMS), which offer low power consumption and fast response. Moreover, each FCA is equipped with local computing hardware, i.e., an LPU, which carries out the necessary baseband processing tasks, such as channel estimation and signal precoding and combining, in a decentralized and parallel fashion, as shown in Fig. 1. Each LPU is connected to the common CPU at the base station (BS) via separate flexible cables, such as coaxial cables. These cables supply power to the FCA and also facilitate control and signal exchange between the LPU and the CPU. The users are each assumed to be equipped with a single fixed-position antenna. In this architecture, the signals radiated by the passive couplers are generated through induced currents excited by their associated active antenna via near-field electromagnetic coupling, without requiring additional RF chains or feeding networks. By reconfiguring only the passive coupler positions within a small local region, the proposed FCA array enhances wireless communication performance while avoiding the movement of active antennas or RF chains, thereby significantly reducing mechanical control complexity and energy consumption.
@@ II-B Channel Model
As illustrated in Fig. 1, we consider a downlink multiuser wireless communication system where an FCA array is deployed at the BS. The BS serves $K$ single-antenna users. The FCA array can be reconfigured by mechanically adjusting the positions of its passive couplers.
@@
where $\mathbf{p}_{m}=[\mathbf{p}_{1,m}^{\mathrm{T}},\mathbf{p}_{2,m}^{\mathrm{T}},\ldots,\mathbf{p}_{N,m}^{\mathrm{T}}]^{\mathrm{T}}\in\mathbb{R}^{2N}$ is the position vector of the couplers in the $m$-th FCA.
@@
where $\mathbf{a}_{\mathrm{C},k,\ell}(\mathbf{p}_{m})$ denotes the coupler steering vector associated with user $k$, path $\ell$, and the $m$-th FCA, and is given by
@@
Based on (3) and (9), the overall multipath channel vector from the FCA array to user $k$ is given by
@@
which depends on coupler positions $\mathbf{p}$ while the active antenna positions are fixed. Collecting the channel vectors of all users yields the following multiuser MIMO channel matrix
@@ II-C Mechanical Beamforming
Let $\widetilde{\mathbf{Z}}\in\mathbb{C}^{M\times M}$ denote the mutual impedance among all active antennas and $\hat{\mathbf{Z}}(\mathbf{p})\in\mathbb{C}^{MN\times MN}$ denote the mutual impedance among all couplers. The overall mutual impedance matrix of the FCA array is thus given by
@@
For practical implementation considerations, we assume that the active antennas are physically isolated by metallic plates. The center-to-center distance between any two active antennas is assumed to be significantly larger than the distance between each active antenna and its associated passive couplers, as well as the distances among the couplers themselves. Consequently, the mutual coupling between ports associated with different FCAs becomes negligible and can be ignored. Under this assumption, the overall mutual impedance matrix exhibits a block-diagonal structure. Accordingly, we define the mutual impedance sub-matrix corresponding to the $m$-th FCA as
@@
Let $X_{n,m}$ denote the load impedance at the port of the $n$-th coupler associated with the $m$-th active antenna. The diagonal load-impedance matrix connected to the $N$ couplers of the $m$-th FCA is given by [9]
@@
Let $s\in\mathbb{C}$ denote the transmit information symbol with zero mean and unit variance. Antenna current vector $\mathbf{i}$ is adopted as the equivalent transmit precoding (beamforming) vector, such that the radiated signal is given by $\mathbf{i}s$. For the $m$-th FCA, define $\mathbf{v}_{\mathrm{C},m}\in\mathbb{C}^{N\times 1}$ as the port-voltage vector across the load impedances at the network side, and define the load-voltage drop along the direction of coupler current vector $\mathbf{i}_{\mathrm{C},m}\in\mathbb{C}^{N\times 1}$. By Kirchhoff’s voltage law [12], we have
@@
is the local precoding/beamforming vector of the $m$-th FCA antenna. We refer to $\mathbf{w}_{m}(\mathbf{p}_{m})$ as the mechanical beamforming vector, since its beamforming weights can be reconfigured through adjusting the coupler positions, $\mathbf{p}_{m}$. Stacking the mechanical beamforming vectors of all FCAs gives the following block-diagonal mechanical beamforming matrix,
@@ II-D Signal Model
For a given coupler-position vector $\mathbf{p}$, we use the active-antenna current vector as the equivalent digital beamforming vector across the $M$ FCAs. Specifically, for user $k\in\mathcal{K}$, we define
@@ III-B MMSE-Based Digital Beamforming
For given coupler positions $\mathbf{p}$, the FCA system reduces to a conventional downlink multiuser MIMO with precoder $\mathbf{F}$ under (48) and effective channel matrix
@@ III-C Coupler Position Optimization Algorithm
Problem (P2) is still nonconvex due to the complicated dependence of $R_{\rm MMSE}(\mathbf{p})$ on $\mathbf{p}$ and the nonconvex spacing constraints in (42d). However, the two constraints (42c) and (42d) are separable across different FCAs, which enables a distributed optimization algorithm.
@@
denote the local feasible set of coupler positions at the $m$-th FCA. Due to the spacing constraints, $\mathcal{S}_{m}$ is in general nonconvex. To apply the SCA framework with convergence guarantees, we construct at each iteration a convex inner approximation of $\mathcal{S}_{m}$. Specifically, for each pair $(n,n^{\prime})$ with $n\neq n^{\prime}$, we define
@@
For problem (P2), the feasible set is separable across the $M$ FCAs through $\{\mathcal{S}_{m}\}_{m=1}^{M}$, while the objective function $R_{\rm MMSE}(\mathbf{p})$ couples all antennas through the effective channels and the MMSE-based digital precoder. Let
@@
Algorithm 1 summarizes the distributed coupler position optimization procedure based on a distributed SCA scheme, where each FCA maximizes a strongly concave local surrogate of the global sum rate. Under standard smoothness assumptions on $J(\mathbf{p})$ and a diminishing step size sequence, the generated sequence $\{\mathbf{p}^{(t)}\}$ yields nondecreasing sum rate $J(\mathbf{p}^{(t)})$, and every limit point of $\{\mathbf{p}^{(t)}\}$ is a stationary solution of problem (P2). The per-iteration computational complexity is dominated by the computation of the MMSE-based precoder and the gradients, and is given by $\mathcal{O}(K^{2}M+K^{3}+KMN+MN^{2})$. Let $T_{\max}$ denote the number of iterations required by the stopping criterion in Algorithm 1. The overall complexity therefore scales as $\mathcal{O}\!\big(T_{\max}(K^{2}M+K^{3}+KMN+MN^{2})\big)$. At each iteration, the CPU broadcasts the local gradients to all LPUs, and each LPU uploads its updated coupler-position vector to the CPU. Hence, the communication overhead scales on the order of $MN$ scalar parameters.